### v3

# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

Abstract.
Regular Expression Denial of Service (ReDoS) is a vulnerability class that has become prominent in recent years. Attackers can weaponize such weaknesses as part of asymmetric cyberattacks that exploit the slow worst-case matching time of regular expression (regex) engines. In the past, problematic regular expressions have led to outages at Cloudflare and Stack Overflow, showing the severity of the problem. While ReDoS has drawn significant research attention, there has been no systematization of knowledge to delineate the state of the art and identify opportunities for further research. In this paper, we describe the existing knowledge on ReDoS. We first provide a systematic literature review, discussing approaches for detecting, preventing, and mitigating ReDoS vulnerabilities. Then, our engineering review surveys the latest regex engines to examine whether and how ReDoS defenses have been realized. Combining our findings, we observe that (1) in the literature, almost no studies evaluate whether and how ReDoS vulnerabilities can be weaponized against real systems, making it difficult to assess their real-world impact; and (2) from an engineering view, many mainstream regex engines now have ReDoS defenses, rendering many threat models obsolete. We conclude with an extensive discussion, highlighting avenues for future work. The open challenges in ReDoS research are to evaluate emerging defenses and support engineers in migrating to defended engines. We also highlight the parallel between performance bugs and asymmetric DoS, and we argue that future work should capitalize more on this similarity and adopt a more systematic view on ReDoS-like vulnerabilities.
Systematization of Knowledge (SoK), Regular Expression Denial of Service (ReDoS), Regex Engines, ReDoS Mitigation Strategies
Both authors contributed equally to this research.
ccs: General and reference Surveys and overviewsccs: Security and privacy Denial-of-service attacksccs: Software and its engineering Software libraries and repositoriesconference: ; ;
1. Introduction
Regular Expression Denial of Service (ReDoS) (Crosby, 2003; Goyvaerts, 2003; Roichman and Weidman, 2009) is a type of denial-of-service (DoS) attack (Gligor, 1983; noa, [n. d.]) that exploits inefficiencies in regular expression (regex) engines. Many regex engines have worst-case exponential backtracking behavior. By crafting input strings that trigger this behavior, attackers can cause systems to consume disproportionate CPU resources, leading to service disruptions. ReDoS threatens real systems: slow regex matches caused outages at Stack Overflow (stackstatus, 2016) and Cloudflare (clo, 2019). ReDoS is the fourth most common vulnerability in the JavaScript npm package ecosystem (Bhuiyan et al., 2023) and one of the fastest-growing vulnerability classes within the same ecosystem (sny, 2019), triggering more than one million pull requests (Section D.5). Consequently, there has been much research on ReDoS, with dozens of papers since 2015 (Section 3). However, as yet there has been no systematization of knowledge about ReDoS, making it challenging to consolidate existing knowledge, identify gaps, and guide future efforts to address this complex and evolving threat.
To address this gap, we conducted a comprehensive literature review of recent papers on ReDoS (Section 3). Our goals were to provide a tutorial on regular expressions and ReDoS, summarize previous work, assess progress in the field, and identify future work opportunities. We analyze 35 papers on ReDoS vulnerabilities from top security conferences, focusing on detection, prevention, and mitigation strategies, to provide a comprehensive overview of the current state of the field. One primary observation from our review was that almost all prior works rely on at least one of two main models of regex engine behavior: the first introduced in 1968 by Thompson (Thompson, 1968), and the second in 1994 by Spencer (Spencer, 1994). While these models have been instrumental in shaping foundational research, they do not reflect modern regex engines, many of which now incorporate defenses against ReDoS attacks. ReDoS research must be situated within real regex engine implementations, motivating us to perform an updated engineering review.
Thus, we conducted an engineering review of modern regex engines to address this disconnect between academic research and practical applications and update the understanding of the capabilities of regex engines (Section 4). This review examines the implementations in the major programming languages to evaluate how they handle ReDoS vulnerabilities, either through new algorithms, resource caps, or other mitigations. By analyzing these engines’ designs, defenses, and real-world adoption, we provide a detailed account of the current state of regex processing and its implications for ReDoS research. This engineering perspective complements the findings of our literature review and highlights the differences between traditional regex engines and contemporary regex engine behavior, offering up-to-date insights for future studies.
As a result of our synthesis of academic knowledge and engineering analysis, we report five findings from this systematization of knowledge. In our review of the academic literature, we identified areas where threat models, ReDoS definitions, and attack evaluations could benefit from further refinement and alignment. In our engineering review of the regex engines, we note that four major engines have used three distinct algorithms to mitigate ReDoS, but these defenses do not always reduce the time complexity to linear. We encourage the research community to explore opportunities to improve conceptual clarity in future ReDoS research and to focus efforts on evaluating and advancing state-of-the-art ReDoS defenses. To support researchers, we propose a framework described in Appendix C, which provides a uniform evaluation environment and also allows comparisons between different regex engines.
Our primary contributions are:
  • We analyze 35 papers on ReDoS vulnerabilities from top security conferences, critiquing their definitions, methodologies, and assumptions. By identifying gaps in the literature, we provide a foundation for more practical and impactful research.
  • We examine the current posture of the major regex engines, highlighting recent improvements, persistent vulnerabilities, and the trade-offs involved in mitigating ReDoS.
  • Based on our findings, we propose guidelines for researchers to align their work with real-world needs and for practitioners to better evaluate and mitigate ReDoS risks.
2. Background and Motivation
In this section, we describe regexes, regex engines, and ReDoS.
2.1. Regexes
2.1.1. Kleene’s regexes (K-regexes)
Regexes are a popular technology for string matching tasks (Friedl, 2006). Regexes are used in two basic operations: recognition (“Does this regex match this string?”) and parsing (“Extract relevant substrings from this match”) (Hopcroft et al., 2007).
Regexes began with Kleene’s (Kleene, 1956) “regular” notation (K-regexes), which specified a language as a set of strings using core constructs of concatenation (.), disjunction (|), and unbounded repetition (*). These operations define whether a given string belongs to the set of strings described by the regex. Formally, a K-regex $R$ is:
$R\to\emptyset\;\Big{|}\;\epsilon\;\Big{|}\;\sigma\;\Big{|}\;R*\;\Big{|}\;R_{1}\mid R_{2}\;\Big{|}\;R_{1}\cdot R_{2}$
where the empty language is $\emptyset$, the empty string is $\epsilon$, and the alphabet is $\Sigma$. A regex’s semantics are defined by its language, the set of strings it describes. The language function $L:R\to 2^{\Sigma^{*}}$ is:
$L(\phi)=\phi$ $L(R\ast)=L(R)\ast$
$L(\epsilon)=\{\epsilon\}$ $L(R_{1}|R_{2})=L(R_{1})\cup L(R_{2})$
$L(\sigma)=\{\sigma\}$ $L(R_{1}{\cdot}R_{2})=L(R_{1}){\cdot}L(R_{2})$
K-regexes can be modeled with nondeterministic or deterministic finite automata (NFA and DFA) (Thompson, 1968; McNaughton and Yamada, 1960). These automata are 5-tuples $\langle Q,q_{0},F,\Sigma,\delta\rangle$, where $Q$ is the set of states, $q_{0}$ is the start state, $F$ is the set of accepting states, $\Sigma$ is the input alphabet, and $\delta$ is the transition function. Each regex operator—character matching, concatenation, repetition, and disjunction—has a corresponding NFA representation, as illustrated in Figure 1. These NFAs can be further converted into DFAs, providing deterministic evaluation at the cost of potentially higher state complexity (Rabin and Scott, 1959).
$q_{0}$$q_{1}$$\sigma$
   $\sigma$
[…]
   $R\ast$
[…]
   $R_{1}\mid R_{2}$
[…]
   $R_{1}\cdot R_{2}$
Figure 1. K-regex operators and NFA equivalents, following the Thompson-McNaughton-Yamada construction (Thompson, 1968; McNaughton and Yamada, 1960).
2.1.2. Extended regexes (E-regexes)
Over time, to simplify the representation, the K-regexes were “extended” (E-regexes) with additional operators and constructs. Syntax sugar notations, like character ranges (e.g., [A-Z] for A|B|...|Z) and one-or-more-repetition (R+), are still regular and can be transformed into equivalent K-regexes. K-regexes were also enhanced (Friedl, 2006; Hazel, 1997) with irregular features such as backreferences and zero-width assertions (Aho, 1980).
Yet, irregular features increase expressive power but exceed the boundaries of regular languages (i.e., they cannot be directly modeled with finite automata), complicating theoretical analysis, implementation, and matching performance. A complete set of regex notation is given in Appendix A.
2.2. Regex Engines
Regexes are supported in all mainstream programming languages (Friedl, 2006) and are part of many other systems, e.g., intrusion detection and prevention systems (Roesch, 1999; Reg, [n. d.]), web application firewalls (Waf, [n. d.]), and logging tools (Buchanan, 2023; Acc, [n. d.]). These systems all implement or integrate a component called a regex engine that provides support for the regex tasks of recognition and parsing. As summarized by Davis (Davis, 2020b), typical regex engine implementations followed one of two algorithms. The first is a depth-first, backtracking search based on a library by Spencer (Spencer, 1994). The second is a breadth-first, “lockstep” search proposed by Thompson (Thompson, 1968)
The characteristics of these engine styles are summarized in Table 1. For the simple case of K-regexes, both can be modeled as NFA simulations. [^1] Implementations of the Thompson-style typically have a direct mapping to an NFA simulation, while implementations of the Spencer-style may be implemented as recursive descent parsers. Here, we treat K-regexes, where both amount to NFA simulations. These approaches differ in their handling of ambiguity (Allauzen et al., 2008)—where multiple transitions in the NFA are possible.
  • Thompson’s algorithm resolves ambiguity by tracking the set of all states in which the automata might be, updating this set using the transition function $\delta$ on each new character.
  • Spencer’s algorithm resolves ambiguity by trying one path, and backtracking to try the other(s) if the first path fails.
Table 1. Tradeoffs of Spencer and Thompson algorithms for regex engines (Cox, 2007; Davis, 2020b). Thompson’s algorithm uses an efficient algorithm over an automata representation to obtain linear time complexity (in input string length), but this formal approach reduces its expressiveness. Spencer gains E-regex support at the cost of exponential match times.
Algorithm Strategy Match time Expressiveness
Thompson’s BFS NFA sim. Linear match Limited E-regex support
Spencer’s Backtracking Exponential match Easy E-regex support
Most of the time (for most regexes, for most inputs, in most application contexts) both approaches run quickly enough, and do not form a bottleneck for the broader application. However, while the Thompson approach guarantees match times that are linear in the input length, the Spencer approach exhibits worst-case polynomial or exponential match times on certain regex-input combinations. Figure 2 gives two examples where Spencer’s algorithm exhibits polynomial and exponential time complexity, respectively. In each case, on an input $w=aa...ab$, the regex engine encounters ambiguity when processing an “a” symbol. At the ultimate mismatch (caused by the final “b”), the backtracking will take $|w|^{2}$ time for example (a) and $2^{|w|}$ time for example (b). This worst-case behavior is colloquially termed “catastrophic backtracking.” Note that the actual cost of the match depends on the length of the input; the term pumping refers to the repetition of the problematic substring of the input (each “a” in this example).
$q_{0}$$q_{1}$$\epsilon$$a$$a$
   (a) NFA for $a{\ast}a{\ast}$
$q_{0}$$q_{1}$$a$ $q_{2}$$\epsilon$ $a$$a$
   (b) NFA for $(aa{\ast}){\ast}$
Figure 2. Example regexes that, in Spencer’s algorithm, exhibit time complexity that is either: (a) polynomial or (b) exponential; in the length of the input string “aa…ab”.
[…]
2.3. Regex Denial of Service (ReDoS)
2.3.1. ReDoS in Principle
This engineering tradeoff—prioritizing expressiveness and accepting poor worst-case performance—has led to a security vulnerability called ReDoS (Crosby, 2003). ReDoS is assigned CWE-1333: Inefficient Regular Expression Complexity (MITRE, 2024). ReDoS is an example of an algorithmic complexity attack (Crosby and Wallach, 2003), characterized by the small cost to create a malicious input and the large cost to process it. Algorithmic complexity attacks are one kind of asymmetric DoS attack (Mantas et al., 2015), where the asymmetry comes from the worst-case performance of the data structures and algorithms used to implement the system.
Following Davis et al.’s (Davis et al., 2021) definition, the threat model for a ReDoS attack has three ingredients:
  1. (1)
    A Backtracking Regex Engine: The regex engine employed for pattern matching in the victim system utilizes a backtracking approach, as discussed in Section 2.2. Many regex engines implemented in programming languages (e.g., JavaScript, Java, Python) have at some point met this criterion (Section 4).
  2. (2)
    A Vulnerable Regex: The victim system uses a super-linear (SL), problematically ambiguous regex; for which the attacker crafts attack strings that take polynomial or exponential time to match in the string length (Wüstholz et al., 2017; Weideman et al., 2016).
  3. (3)
    An Adversary-Controlled Input: The attacker supplies inputs that result in attack strings reaching the vulnerable regex and triggering its worst-case time complexity.
Given these three ingredients, if the system has insufficient mitigations (safeguards) in place, then the attacker can cause the victim system to consume excessive computational resources (CPU time). The attacker may need to tailor the number of pumps based on the length of typical input (to avoid anomaly detection) or the maximum length allowed (to avoid truncation).
Figure 3. Distribution of OWASP Top-10 and ReDoS CVEs per ecosystem, showing average prevalence per category. Data cover 2014-2023. For full methodology, see Appendix B.
2.3.2. ReDoS in Practice
In recent years, ReDoS has become a common security concern. Two major web services had outages that resulted from slow regex behavior: Stack Overflow in 2016 (stackstatus, 2016) and Cloudflare in 2019 (clo, 2019). While these outages did not involve an adversary (i.e., they were reliability issues, not security issues), they made service providers aware of the risk of ReDoS. Since 2020, the security company Snyk has observed a jump in reported ReDoS vulnerabilities, especially in the JavaScript package ecosystem npm (sny, 2019). Grey literature resources such as GitHub’s security team give advice for repairing ReDoS vulnerabilities (sny, [n. d.]; Backhouse, 2023; Davis, 2020a; reg, [n. d.]).
For a sense of the formal disclosure of ReDoS vulnerabilities, we examined the National Vulnerability Database for CVE records pertaining to ReDoS. Figure 3 shows the result. Following the keyword method of Hassan et al. (Hassan et al., 2023), since 2014 $\sim$400 ReDoS CVEs have been disclosed in software from six major programming language ecosystems. As a reference point, we compared these results to CVEs related to the “OWASP Top-10”. Averaging over 2014-2023, the rate of ReDoS exceeds four of the OWASP Top-10 vulnerability types, though it is far below the Top-5.
3. Academic Research on ReDoS
Contemporaneous with the increasing industry recognition of ReDoS, ReDoS has become a common subject of cybersecurity research. However, there has been no systematization of the resulting knowledge. We therefore conducted a systematic literature review of ReDoS papers published between 2015 and 2024 at leading conferences in computer security (S&P, USENIX, CCS, NDSS), software engineering (ICSE, FSE, ASE), and programming languages (PLDI, POPL, OOPSLA). Following the structured methodology outlined by Schloegel et al. (Schloegel et al., 2024), we identified relevant papers in conference proceedings via an initial keyword search conducted in June 2024, using the terms ReDoS, denial-of-service, regex, regular expression, regular expression analysis, and regular expression execution. This yielded 93 papers, which were further filtered via the following criteria:
  • Inclusion Criteria: Papers that primarily address ReDoS, propose new attacks or defenses against ReDoS, or extensively study existing ReDoS attacks or defenses.
  • Exclusion Criteria: Papers that focus on general denial of service or those that discuss regular expressions, without specifically addressing ReDoS.
To ensure informed inclusion and exclusion decisions, we reviewed the abstract, introduction, and evaluation sections of each paper. We identified 35 papers that met our inclusion criteria and systematically analyzed them along four key dimensions:
  1. (1)
    Research Directions: Captures the primary focus of each study, which typically falls into one of three categories:
    • Detection: Identifying and analyzing vulnerabilities.
    • Prevention: Addressing root causes by improving regex engines or limiting unsafe constructs.
    • Mitigation: Reducing attack impact through runtime defenses or repair mechanisms.
  2. (2)
    Threat Models: Examines the assumptions about attacker capabilities (e.g., control over input) and system behavior (e.g., rate-limiting or regex engine design), highlighting the practical applicability of each study.
  3. (3)
    Vulnerability Definitions: Focuses on how papers describe what constitutes a ReDoS vulnerability or a vulnerable regex, e.g., having exponential matching time.
  4. (4)
    Evaluation Methods: Assesses the languages and regex engines used for testing, providing insights into the platforms where these approaches are validated.
Additionally, we reviewed other closely related works that, while not meeting our strict inclusion criteria, are still highly relevant to ReDoS. For instance, papers such as (Cox, 2007; Holík et al., 2023; Smith et al., 2008) propose improved regex matching algorithms. Although mitigating ReDoS was not their primary focus, these works provide valuable insights and serve as foundational building blocks for ReDoS defenses, making them significant for academics and researchers in the field.
We begin by discussing the main existing research directions (Section 3.1), distinguishing between detection, prevention, mitigation, and other related works. We then explore the assumptions made about the attackers’ capabilities or the underlying systems (Section 3.2), along with the definitions of ReDoS vulnerabilities or successful attacks provided in prior works (Section 3.3). Finally, we discuss how the studied papers were evaluated (Section 3.4).
{forest}
[…]
Figure 4. Systematization of ReDoS research papers categorized into detection, prevention, mitigation, and related studies, with subcategories and key references.
3.1. Main Existing Research Directions
Figure 4 provides an overview of the main research directions in the study of ReDoS vulnerabilities, categorized into detection, prevention, mitigation, and complementary studies. Detection focuses on identifying ReDoS vulnerabilities in regular expressions and understanding their exploitability. Prevention aims to address the root causes of ReDoS vulnerabilities by improving regex engines or restricting vulnerable constructs. Mitigation, on the other hand, focuses on minimizing the impact of ReDoS attacks by deploying runtime safeguards. We discuss these categories in detail below.
3.1.1. Detection
Detection is a critical area of research in ReDoS, with studies focusing on identifying vulnerable regexes across various ecosystems. Nine papers analyze open-source libraries, such as npm (Bhuiyan et al., 2023; Staicu and Pradel, 2018), pip (Davis et al., 2018a; Wang et al., 2023), and Maven (Wang et al., 2023; Li et al., 2021), while seven focus on testing live web applications (Staicu and Pradel, 2018; Barlas et al., 2022). Additionally, 27 papers investigate ReDoS vulnerabilities using curated regex datasets (Turoňová et al., 2022; Liu et al., 2021; Siddiq et al., 2024a, b). We categorize detection methods into five key approaches: Heuristic-Based Detection (three papers), Automata-Based Detection (seven papers), Dynamic Detection (four papers), Hybrid Detection (five papers), and Machine Learning-Based Detection (one paper). Heuristic-based techniques rely on simple pattern rules to flag problematic regexes, offering speed but often at the cost of high false positive rates. Automata-based approaches, which model regexes using finite-state machines, provide formal accuracy but are computationally expensive. Dynamic detection evaluates regex behavior on specific inputs, revealing real-world vulnerabilities but requiring significant runtime resources. Hybrid methods combine static and dynamic techniques to balance precision and efficiency, while emerging machine learning approaches use regex datasets to predict vulnerabilities, opening new avenues but facing challenges like adversarial examples and limited training data. This categorization highlights the trade-offs in speed, accuracy, and false positive/negative rates among different techniques, offering a structured view of existing detection strategies and their potential for future enhancement.
Heuristic-Based Detection (3 studies) There are many approaches that employ simple pattern-matching techniques to identify potentially vulnerable regular expressions. Studies by Kluban et al. (Kluban et al., 2024, 2022) and Davis et al. (Davis et al., 2017) utilize tools like regexploit (noa, 2024b), redos-detector (jen, 2024), and safe-regex (noa, 2024a) to identify these vulnerabilities by searching for constructs such as infinite repeats (a*b*a*), branches ($a|b$), or nested quantifiers (*, +, ?, {n,m}), which are often associated with ReDoS vulnerabilities. While these methods are fast, enabling the rapid analysis of large numbers of regular expressions, they typically focus on syntax rather than semantics, leading to potential accuracy issues. For instance, Parolini et al. (Parolini and Miné, 2023) demonstrated that while tools like regexploit achieved low false positive and negative rates, others like safe-regex and redos-detector exhibited significantly higher false positive rates. This highlights the primary limitation of heuristic-based approaches, which often require further analysis to confirm vulnerabilities.
Automata-based Static Analysis (6 studies) Automata-based static analysis is a powerful approach for detecting problematic regular expressions without requiring program execution. This method has been explored in several studies (Rathnayake and Thielecke, 2017; Kirrage et al., 2013; Parolini and Miné, 2023; Turoňová et al., 2022; Su et al., 2024; Weideman et al., 2016), which employ automata models such as NFAs, DFAs, and enhanced automata to evaluate regex semantics and detect vulnerabilities that lead to ReDoS attacks.
NFA-based methods, such as those by Kirrage et al. (Kirrage et al., 2013) and Rathnayake et al. (Rathnayake and Thielecke, 2017), construct NFAs to analyze patterns like (prefix, pumpable string, suffix) and identify vulnerabilities caused by exponential backtracking. Building on this, Weideman et al. (Weideman et al., 2016) enhance NFAs with prioritization (pNFA) to reduce redundant state exploration, providing a more precise analysis of backtracking vulnerabilities.
DFA-based approaches, as demonstrated by Turoňová et al. (Turoňová et al., 2022) and Shen et al. (Su et al., 2024), rely on deterministic automata simulations to evaluate vulnerabilities in non-backtracking engines. By incorporating specialized counting mechanisms, these approaches effectively detect problematic patterns, such as bounded quantifiers and repetitive constructs, while optimizing resource usage. This focus on determinism allows for efficient analysis of regexes without the ambiguity associated with backtracking.
Enhanced automata models offer a novel perspective on regex mathing. Parolini et al. (Parolini and Miné, 2023) introduce tree semantics, extending beyond traditional NFA and DFA approaches by capturing the structural interactions of regex engines. This enables the detection of vulnerabilities that are difficult to represent in standard automata.
While these methods excel in early detection, they face challenges such as false positives from over-approximations, false negatives from under-approximations, and high computational complexity for analyzing intricate regex structures. Despite these trade-offs, automata-based static analysis remains a foundational tool for addressing ReDoS vulnerabilities.
Dynamic Analysis (4 studies) Dynamic approaches (Shen et al., 2018; McLaughlin et al., 2022; Barlas et al., 2022; Petsios et al., 2017) to detecting ReDoS vulnerabilities involve executing regular expressions with various inputs and analyzing their runtime behavior to identify potential issues. For instance, Shen et al. (Shen et al., 2018) developed ReScue, a tool that uses an evolutionary genetic algorithm to generate time-consuming strings capable of triggering ReDoS vulnerabilities. Similarly, Barlas et al. (Barlas et al., 2022) and McLaughlin et al. (McLaughlin et al., 2022) utilized dynamic analysis with fuzzing techniques, while Petsios et al. (Petsios et al., 2017) introduced SlowFuzz, which uses automated fuzzing to find inputs causing worst-case algorithmic behavior.
Hybrid Approaches (6 studies) Existing static analysis methods for detecting ReDoS vulnerabilities often face a trade-off between precision and recall. For example, Davis et al. (Davis et al., 2018a) and Shen et al. (Shen et al., 2018) reported low F-1 scores of 44.94% and 3.57%, respectively (Li et al., 2021). To address these limitations, dynamic validation is frequently employed alongside static analysis, resulting in hybrid approaches. Tools such as NFAA (Wüstholz et al., 2017), Revealer (Liu et al., 2021), ReDoSHunter (Li et al., 2021), Badger (Noller et al., 2018), and RENGAR (Wang et al., 2023) combine static and dynamic methods to mitigate false positives. For instance, hybrid approaches like ReDoSHunter integrate static and dynamic techniques to improve detection accuracy, while tools such as NFAA refine dynamic testing with insights from static analysis (Wüstholz et al., 2017). These methods effectively balance the strengths of both static and dynamic detection, reducing the likelihood of inaccurate results.
Machine Learning Approaches (1 studies) Machine learning techniques are also gaining traction in ReDoS detection. Unlike hybrid methods, which explicitly integrate static and dynamic analysis, machine learning approaches aim to predict vulnerabilities and detect anomalies based on patterns in data. For example, Demoulin et al. (Demoulin et al., 2019) introduced a semi-supervised learning model to analyze resource utilization and identify potential vulnerabilities. This method triggers immediate alerts and provides mitigation strategies by detecting anomalies in resource consumption. While machine learning approaches offer promise, they face challenges such as reliance on labeled data, vulnerability to adversarial inputs, and difficulty in addressing context-specific patterns that may not generalize well across diverse regex use cases.
While dynamic and hybrid techniques are more precise, they tend to be slow. Parolini et al. (Parolini and Miné, 2023) show that ReScue (Shen et al., 2018) and Rexploiter (Wüstholz et al., 2017) are 2055x and 57x slower than RXXR2 (Rathnayake and Thielecke, 2017). Thus, despite their strenghts, these methods exhibit slow execution time and incomplete coverage.
3.1.2. Prevention
Prevention strategies focus on addressing the root causes of ReDoS vulnerabilities, either by improving regex matching algorithms or by repairing vulnerable regex patterns. These approaches aim to prevent ReDoS attacks entirely by designing systems that eliminate worst-case performance scenarios, thus avoiding the need for runtime detection or mitigation.
Better Regex Matching Algorithms (3 studies) Researchers have proposed developing new regex matching engines to address ReDoS vulnerabilities at a fundamental level. Davis et al. (Davis et al., 2021; Davis, 2019) introduced memoization-based optimization to speed up regex matching by caching intermediate results, significantly reducing redundant computations. Turoňová et al. (Turoňová et al., 2020, 2022) proposed novel counting automata matching algorithms that further optimize counter management during matching to efficiently handle regexes with bounded repetitions. These new engines often employ finite-state machine-based approaches instead of backtracking, inherently avoiding the exponential time complexity that leads to ReDoS vulnerabilities. Additionally, some methods impose restrictions on regex features to ensure faster and safer processing.
Automatic Repair (4 studies) Automatic repair techniques aim to proactively address vulnerabilities in regular expressions to mitigate ReDoS attacks. Several studies have explored programmatic frameworks to detect and repair problematic regex patterns. For example, Li et al. (Li et al., 2020) developed FlashRegex, a programming-by-example (PBE) tool that repairs or synthesizes regex from provided matching examples. However, FlashRegex struggles with complex regex features, such as lookarounds and backreferences (Li et al., 2022). To address these limitations, RegexScalpel (Li et al., 2022) employs a localize-and-fix approach, which identifies fine-grained vulnerabilities and applies predefined repair strategies to improve recall and precision. In a complementary effort, Hassan et al. (Hassan et al., 2023) focus on identifying specific regex structures or “anti-patterns” that are prone to ReDoS vulnerabilities. Their method analyzes regex patterns to uncover and repair potential vulnerabilities early in the development process, avoiding the need for runtime execution. This approach complements programmatic frameworks by systematically identifying and addressing problematic constructs.
Other tools, such as Chida et al. (Chida and Terauchi, 2022), also leverage PBE algorithms but face challenges related to dependency on user-provided examples, which can result in incomplete repairs (Li et al., 2021; Wang et al., 2023). These methods, while effective for certain regex patterns, often struggle with advanced or highly customized regex constructs, underscoring the need for more comprehensive solutions capable of addressing complex vulnerabilities. Moreover, improving the generalization and robustness of automatic repair frameworks remains a key area of ongoing research.
Prevention techniques address ReDoS vulnerabilities by improving the efficiency and resilience of regex engines or by repairing vulnerable regex patterns. While they significantly reduce the risk of attacks, these methods require careful configuration to ensure compatibility with complex patterns and real-world scenarios.
3.1.3. Mitigation
Mitigation strategies aim to detect ReDoS attacks in live systems and mitigate their impact in real-time. These approaches focus on monitoring runtime behavior to identify anomalies in resource consumption, such as excessive memory usage or prolonged processing times, and applying defensive measures to minimize the effects of ongoing attacks. Unlike prevention, which aims to eliminate vulnerabilities before execution, mitigation provides a dynamic response to attacks as they occur.
Anomaly Detection (2 studies) Anomaly detection techniques are central to mitigation strategies, leveraging real-time monitoring to identify unusual patterns indicative of ReDoS attacks. Machine learning (ML)-based approaches have proven effective in this domain (Bai et al., 2021; Tandon et al., 2023). For example, Tandon et al. (Tandon et al., 2023) employ the Elliptic Envelope model to analyze runtime features such as request time, memory usage, and CPU cycles, enabling the system to detect deviations from normal behavior. Bai et al. (Bai et al., 2021) introduce a feedback loop mechanism that continuously adjusts defensive measures based on real-time performance metrics, providing robust protection against zero-day ReDoS attacks.
While promising, ML-based anomaly detection methods are not without challenges. For instance, Tandon et al. (Tandon et al., 2023) report a false positive rate of 1.3% and a false negative rate of 0.4%, which could still impact practical deployment. Furthermore, these systems can be vulnerable to adversarial attacks, where carefully crafted inputs are designed to bypass detection mechanisms (Bai et al., 2021). These limitations underscore the need for more resilient and adaptive approaches to anomaly detection, ensuring robustness against both known and emerging threats.
Limiting Resource Consumption (1 study) To mitigate ReDoS attacks, Davis et al. (Davis et al., 2018b) proposed using first-class timeouts as a defense against Event Handler Poisoning (EHP) in Node.js. Their prototype, Node.cure, effectively limits processing time with minimal performance overhead. Similarly, runtime environments like PHP, Perl, and .NET have adopted resource-limiting mechanisms to curb backtracking and excessive processing. In PHP, the pcre.backtrack_limit and pcre.recursion_limit directives control the number of backtracks and recursion depth during regex matching (PHP Documentation Team, 2025). Perl’s PCRE2 library allows setting evaluation limits via the eval function, with a default recursion depth limit of 10M. In .NET, the RegexOptions.MatchTimeout property enables developers to specify a maximum regex processing duration. While these strategies effectively cap resources to reduce ReDoS risks, they are often too coarse-grained, making it challenging to balance performance and avoid false positives in complex regex patterns or large datasets.
3.1.4. Other Related Work
Beyond approaches for addressing ReDoS directly, several complementary studies provide valuable insights into datasets, real-world analyses, and novel frameworks.
Dataset Studies Bhuiyan et al. (Bhuiyan et al., 2023) introduced SecBenchJS, a benchmark dataset designed to evaluate the security of JavaScript applications, including regex vulnerabilities. Such datasets enable standardized testing and improve ReDoS detection tools.
Empirical Studies Staicu et al. (Staicu and Pradel, 2018) analyzed 2,846 popular websites, finding 339 vulnerable to ReDoS attacks, highlighting the risks of server-side JavaScript. Davis et al. (Davis et al., 2019) studied regex portability across eight programming languages, uncovering semantic and performance differences in 25% of regexes. These studies underscore the need for safer regex practices and consistent regex engine implementations.
LLM-Based Studies Siddiq et al. (Siddiq et al., 2024b) proposed Re(gEx—DoS)Eval, a framework for evaluating LLM-generated regexes using metrics like pass@k and vulnerable@k, demonstrating its use on T5, Phi, and GPT-3. In another study, Siddiq et al. (Siddiq et al., 2024a) categorized vulnerable regex patterns and analyzed developer discussions, finding that GPT-3.5 often generates regexes with polynomial vulnerabilities.
These studies provide benchmarks, empirical data, and frameworks that complement prevention, detection, and mitigation strategies, paving the way for robust ReDoS solutions.
3.2. Threat Models
Table 2. Overview of threat models considered in the reviewed studies. Most works consider regexes in isolation or only within a library; only 7/35 of these works consider reachability at the application level. Beyond this, ReDoS research makes strong assumptions about the attacker’s capabilities (or assumes a poorly defended system).
Paper Code Fragment Library Application Arbitrary Number of Requests Su et al. (Su et al., 2024) 100kB Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) 1.5K-9K Liu et al. (Liu et al., 2021) 128 Staicu et al. (Staicu and Pradel, 2018) 82K Davis et al. (Davis et al., 2018a) 100K-1M Siddiq et al. (Siddiq et al., 2024a) Shen et al. (Shen et al., 2018) 128 Barlas et al. (Barlas et al., 2022) Siddiq et al. (Siddiq et al., 2024b) Davis et al. (Davis et al., 2019) Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) Noller et al. (Noller et al., 2018) Wang et al. (Wang et al., 2023) Rathnayake et al. (Rathnayake and Thielecke, 2017) McLaughlin et al. (McLaughlin et al., 2022) 1M Li et al. (Li et al., 2021) Parolini et al. (Parolini and Miné, 2023) 128 Kirrage et al. (Kirrage et al., 2013) Wüstholz et al. (Wüstholz et al., 2017) 140K Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) 50K Li et al. (Li et al., 2022) 1M Davis et al. (Davis, 2019) Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) Davis et al. (Davis et al., 2021) 10K-100K Tandon et al. (Tandon et al., 2023) Weideman et al. (Weideman et al., 2016) Turoňová et al. (Turoňová et al., 2020) 500K-50M
A realistic threat model is essential in any security study, as it provides a framework for assessing the potential findings. That is, it enables the identification of system’s weaknesses, i.e., vulnerabilities, that adversaries could exploit in real-world environments. A strong threat model might lead to unrealistic findings, assuming too powerful attackers. In  Table 2, we outline the granularity at which different threat models are defined, in the analyzed papers. While most studies examine the exploitability of regular expressions in isolation, some do extend their analysis to the surrounding context, such as within library code. However, only a few studies define threat models at the application level. Threat models defined at code fragment level can lead to false positives—vulnerabilities that may be exploitable in isolated regular expressions but remain unreachable in actual code.
Among the papers we reviewed, only 10 (28%) explicitly define a threat model, while the rest implicitly assume one. We identify four assumptions in the existing threat models:
Attacker Controls Input (35 studies): The assumption that an attacker can control inputs to a vulnerable ReDoS (Regular Expression Denial of Service) location is crucial for exploitability, as highlighted in various studies (Barlas et al., 2022; Hassan et al., 2023). In real-world software, including web applications and systems, inputs often originate from untrusted sources, thereby increasing the potential for exploitation (Petsios et al., 2017). However, whether these inputs can reach the vulnerable regular expression depends on the surrounding code. All the reviewed papers, whether their threat models focus on the application (Staicu and Pradel, 2018; Barlas et al., 2022; Wüstholz et al., 2017; Demoulin et al., 2019), library (Bhuiyan et al., 2023; Wang et al., 2023; McLaughlin et al., 2022; Li et al., 2021), or code fragment level (Davis et al., 2017; Turoňová et al., 2022; Liu et al., 2021), assume that the attacker can manipulate the input to exploit the vulnerability in the regular expression. However, this assumption may not always be applicable. For instance, consider a library that only reads and parses a configuration file on a server, which is not controlled by an attacker. In such cases, the likelihood of an attacker being able to influence the input and exploit the ReDoS vulnerability is significantly reduced, if not eliminated.
Attacker’s Input Reaches Super-Linear Regex (34 studies): This condition necessitates that the attacker’s input be matched against a specific type of regular expression. While all the papers considering threat models at the library and code fragment levels implicitly assume this condition, it is a significant assumption. With the exception of Barlas et al. (Barlas et al., 2022), most studies presume this might be achievable in practice. In reality, achieving this could require the attacker to have access to detailed information about the server-side logic, API documentation, or the internal workings of the web service. Staicu et al. (Staicu and Pradel, 2018) incorporate web server implementations into their threat model and examine whether the input can trigger the vulnerability. However, this fingerprinting approach can be noisy, leading to both false positives and false negatives.
Use of Backtracking or Slow Regex Engine (33 studies): This condition assumes the attacker has in-depth knowledge of the specific regex engine employed by the server. On one hand, it might be feasible to remotely determine the type of engine in use (e.g., PCRE, RE2), but pinpointing the exact version with its specific weaknesses can be quite challenging. On the other hand, assuming a slow or outdated engine might lead to inaccurate results, making the attack scenario unrealistic in many real-world situations. Except for Barlas et al. (Barlas et al., 2022) and Turoňová et al. (Turoňová et al., 2022), all the studies make this assumption as part of their threat model.
Arbitrary Request Assumption (32 studies): This condition represents a strong assumption that the attacker can repeatedly send requests with large input lengths without interference from other security mechanisms. Most studies, with the exceptions of (Shen et al., 2018; Liu et al., 2021; Parolini and Miné, 2023), either disregard restrictions on request size and frequency or necessitate large inputs to trigger vulnerabilities. For context, the maximum size of HTTP header requests and responses in nginx and Apache Tomcat (apa, [n. d.]; ngi, [n. d.]) is 8K, which is significantly smaller than the input lengths assumed in most works. For instance, Davis et al. (Davis et al., 2018a) require input lengths of 100K-1M to trigger a 10-second slowdown, which would be impractical in real-world system settings. Furthermore, most studies fail to consider potential mitigation strategies employed by real-world systems, such as rate limiting(her, [n. d.]), intrusion detection systems, or firewalls, which can significantly impede attacks. In realistic scenarios, attackers might face various system-imposed limitations, including bandwidth constraints or restrictions on the number of requests(her, [n. d.]) they can send. As a result, the majority of studies, except for (Shen et al., 2018; Liu et al., 2021; Su et al., 2024), assume in their threat models that attackers can perform requests unhindered, which may not reflect real-world conditions.
While the inclusion of a realistic threat model is essential for evaluating security findings accurately, many existing ReDoS studies make unrealistic, strong assumptions about attacker capabilities and system configurations.
3.3. ReDoS Definitions
Table 3. Summary of ReDoS vulnerability definitions in reviewed studies. The literature uses inconsistent criteria, with over half (19/29) defining ReDoS based on slowdown.
Paper Slowdown Number of Instructions Heuristics Su et al. (Su et al., 2024) ¿10000 Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) 10s, 100s Liu et al. (Liu et al., 2021) $10^{5}$ Staicu et al. (Staicu and Pradel, 2018) 5s Davis et al. (Davis et al., 2018a) 10s Siddiq et al. (Siddiq et al., 2024a) 1s Shen et al. (Shen et al., 2018) $10^{8}$ Barlas et al. (Barlas et al., 2022) 1s Siddiq et al. (Siddiq et al., 2024b) 1s $10^{8}$ Davis et al. (Davis et al., 2019) 10s Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) 2s Wang et al. (Wang et al., 2023) $10^{5}$ McLaughlin et al. (McLaughlin et al., 2022) 10s Li et al. (Li et al., 2021) 1s Parolini et al. (Parolini and Miné, 2023) $10^{10}$ Wüstholz et al. (Wüstholz et al., 2017) Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) $\mu+3\sigma$ Li et al. (Li et al., 2022) 10s Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) 1s Davis et al. (Davis et al., 2021) Turoňová et al. (Turoňová et al., 2020) 10s, 100s
Most of the considered papers define a ReDoS vulnerability in terms of the slowdown that can be caused directly, with a limited number of characters. This variation is summarized in Table 3. Notably, there is no consensus on what constitutes a vulnerability. For instance, some studies establish thresholds for the number of allowed operations during matching that differ significantly. Liu et al. (Liu et al., 2021) and Wang et al. (Wang et al., 2023) use a threshold of $10^{5}$ instructions, while Shen et al. (Shen et al., 2018) and Siddiq et al. (Siddiq et al., 2024b) consider $10^{8}$ instructions, and Parolini et al. (Parolini and Miné, 2023) set it at $10^{10}$ instructions. Adding to the confusion, Sung et al. (Sung et al., 2022) define vulnerabilities at $10^{6}$ and $10^{7}$ instructions. It remains unclear why these values are deemed appropriate, as heavy CPU loads do not always indicate the presence of a vulnerability.
Moreover, the inconsistency extends to the measurement of introduced slowdowns. Various papers use different metrics to quantify slowdowns, further complicating the landscape and hindering comparability across studies. Turoňová et al. (Turoňová et al., 2022) measure slowdowns in $10s$ and $100s$, Staicu et al. (Staicu and Pradel, 2018) in $5s$, Davis et al. (Davis et al., 2018a, 2019) in $10s$, Siddiq et al. (Siddiq et al., 2024a), Li et al. (Li et al., 2021) and Barlas et al. (Barlas et al., 2022) in $1s$, Bhuiyan et al. (Bhuiyan et al., 2023) in $2s$. Meanwhile, McLaughlin et al. (McLaughlin et al., 2022) use $10s$, and Bai et al. (Bai et al., 2021) define a problematic slowdown as the average response time ($\mu$) plus three times the standard deviation ($3\sigma$), to account for significant deviations from the norm. These excessively high thresholds likely miss many potential vulnerabilities that could cause significant issues under real-world conditions. In our experiments, we saw that under one-second slowdowns can be weaponized against a target server, with modest attacker resources (Bhuiyan and Staicu, 2022). Other papers rely on heuristics to define a vulnerability. For example, Davis et al. (Davis et al., 2017) utilize safe-regex (noa, 2024a), which uses static analysis rules to determine whether a regex is vulnerable. These tools are unsound and report both false positives and negatives.
The lack of consensus on vulnerability thresholds, slowdown metrics, and measurement methodologies highlights significant inconsistencies in ReDoS research, underscoring the need for standardized definitions and evaluation criteria.
3.4. Evaluations
Table 4. Summary of evaluation methods used in the reviewed studies. Most works (33/34) measure the success of a ReDoS attack from the perspective of the attacker, rather than from its impact on benign users (Quality of Service degradation). Over half of the works (25/34) are evaluated on a single language’s regex engine.
Paper Attack Simulation Measure QoS Degradation JavaScript Java Python Ruby PHP Go Perl Rust C# Su et al. (Su et al., 2024) Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) Liu et al. (Liu et al., 2021) Staicu et al. (Staicu and Pradel, 2018) Davis et al. (Davis et al., 2018a) Siddiq et al. (Siddiq et al., 2024a) Shen et al. (Shen et al., 2018) Barlas et al. (Barlas et al., 2022) Siddiq et al. (Siddiq et al., 2024b) Davis et al. (Davis et al., 2019) Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) Noller et al. (Noller et al., 2018) Wang et al. (Wang et al., 2023) Rathnayake et al. (Rathnayake and Thielecke, 2017) McLaughlin et al. (McLaughlin et al., 2022) Li et al. (Li et al., 2021) Parolini et al. (Parolini and Miné, 2023) Kirrage et al. (Kirrage et al., 2013) Wüstholz et al. (Wüstholz et al., 2017) Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) Li et al. (Li et al., 2022) Davis et al. (Davis, 2019) Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) Tandon et al. (Tandon et al., 2023) Weideman et al. (Weideman et al., 2016) Turoňová et al. (Turoňová et al., 2020)
We carefully studied the evaluations of the papers in our dataset, as shown in Table 4. Our analysis reveals two key aspects regarding ReDoS research:
Language: Regular expression matching is comparatively slow in languages like JavaScript, Java, Python, and Ruby (Davis et al., 2019), leading to a significant amount of research focusing on detecting and defending ReDoS attacks in these languages. Reflecting this emphasis, 19 (54%) papers address JavaScript, 18 (51%) address Java, and 9 (25%) address Python, with JavaScript being particularly prominent.
Evaluation Methodology: Out of the reviewed studies, only 7 (18%), such as Demoulin et al. (Demoulin et al., 2019), Tandon et al. (Tandon et al., 2023), and Staicu et al. (Staicu and Pradel, 2018), performed an end-to-end simulation to evaluate their proposed attack or defense methodologies. Moreover, only Bai et al. (Bai et al., 2021) measured the system’s performance degradation during an attack as perceived by benign users. The remaining 29 studies (82%) rely solely on theoretical or localized analysis, which might not capture real-world attack scenarios and defenses accurately, potentially overestimating the prevalence of ReDoS vulnerabilities.
The lack of comprehensive end-to-end evaluations in current research hampers the ability to accurately gauge the real-world effectiveness and impact of ReDoS attacks and defenses.
4. ReDoS Mitigations in Regex Engines
In Section 3, we discuss many papers that measure ReDoS or propose defenses. They often assume that the matching of a regex engine can be modeled as a backtracking search (Section 3.2). However, it is unclear whether this assumption holds in practice, particularly in the context of modern regex engines.
In this section, we examine the regex engines of major programming languages to answer two questions: (1) Do the measurements and assumptions in the academic literature accurately reflect the latest versions of these engines? and (2) What is the impact and adoption of ReDoS defenses proposed by researchers in real-world regex engines?
We answer these questions in two ways. First, a systematic analysis was conducted on regex engines in nine popular programming languages (Section 4.1). This analysis drew upon first-party sources such as source code, issue trackers, language documentation, blog posts from regex engine maintainers, and CVE reports. We also examined academic literature that describes regex engines. Second, it was measured whether super-linear regexes improve performance in the latest versions of these engines (Section 4.2).
4.1. Regex Engine Analysis
In this section, we begin by outlining the selection criteria used to choose the regex engines for our analysis. Then, for each engine, we examine the type and details of the ReDoS defense(s) implemented (if any), the update context of the defense (including the year of introduction), and the required lines of code (LOC) to implement it. Additionally, we assess whether the defense is enabled by default or requires developer intervention. We summarize our results on evaluating ReDoS defenses in modern regex engines in Table 5.
Table 5. Summary of ReDoS defenses in the implementations of major programming languages.
Language (Implementation) Nature of the Defense Deployment Year LOC to Implement On by Default? Current ReDoS-Safeness JavaScript (Node.js—V8 (Nod, [n. d.]; V8J, [n. d.])) A Non-backtracking Engine: Thompson-style (Thompson, 1968) 2020 1558 (10.6%) Partial Ruby (MRI/CRuby (Rub, [n. d.])) A Non-backtracking Engine: Memoization (Davis et al., 2021), Resource Capping: Time-based 2022 1100 (4.7%) Partial C# (.NET (Wha, [n. d.])) Resource Capping: Time-based; Brzozowski Derivatives Engine (Moseley et al., 2023) 2012, 2022 5417 (34.7%) Partial Java (OpenJDK (Ope, [n. d.])) Caching 2016 1712 (35.5%) Partial PHP (Zend Engine (PHP, [n. d.]; php, [n. d.])) Resource Capping: Counters Partial Perl (perl5 (Per, [n. d.])) Caching, Resource Capping: Counters Partial Rust (rustc (rus, [n. d.])) A Non-backtracking Engine: Thompson-style Safe Go (gc (Int, [n. d.])) A Non-backtracking Engine: Thompson-style Safe Python (CPython (Pyt, [n. d.])) Not Safe
Selection Criteria. Many modern programming languages have multiple implementations, runtime environments, engines, interpreters, or compilers that might use different regex engines. For instance, JavaScript has multiple engines for executing the source code, which can be used in different environments, such as V8 in Node.js and SpiderMonkey in Mozilla Firefox. To address this complexity, we applied the following criteria when selecting a specific regex engine for languages with multiple options:
  1. (1)
    We prioritized engines that are widely used in production environments and are likely to be targets of ReDoS attacks. This criterion ensures our analysis reflects real-world impact and relevance. As an illustrative example, Node.js is used as web server by 3.4% of all websites, and remarkably, it is used by 98.8% of all websites whose server-side programming language is known to be JavaScript (Usa, [n. d.]; Dis, [n. d.]). This makes V8 the most pertinent choice for our analysis. We focus on reference implementations for languages where comprehensive usage data is not readily available. Such implementations are typically mentioned in official documentation of programming languages and serve as the standard for the language (e.g., CPython for Python).
  2. (2)
    Additionally, we only consider implementations with publicly-accessible source code. This allows for a thorough examination of the implementation details and recent modifications to the engine. For instance, in the case of Java, we selected OpenJDK, an open-source platform, over Oracle’s JDK.
An overview of our regex engine selections for this study and their rationale can be found in Section D.1.
4.1.1. JavaScript (Node.js—V8)
Original Algorithm: The main JavaScript engine, V8, previously used a backtracking Spencer-style algorithm. It was highly optimized for common-case regexes (Gruber, 2017). It uses an explicit automaton representation, optimizing it and then rendering it as native machine code for execution (Corry et al., [n. d.]).
ReDoS Defense: In 2020, responding to ReDoS concerns, the V8 developers implemented an extra Thompson-style non-backtracking regex engine (Bidlingmaier, 2021). The new engine guarantees linear-time complexity for all supported regexes, but as a trade-off, it does not support E-regex features. For regexes that are not supported, the original engine serves as a fallback. This update was introduced in V8 version 8.8 (2020), but in Node.js, it remains an experimental feature that is not enabled by default and must be activated via a feature flag. Thus, some Node.js applications may not yet benefit from this feature. All E-regex features still use the backtracking engine.
Engineering Effort: The new engine is 1558 LOC. The original was 14741 LOC (excluding platform-specific assembly code). The effort was $\sim$10of the original.
4.1.2. Ruby (MRI/CRuby)
Original Algorithm: Ruby’s default regex engine is based on Onigmo (k.t, 2024), a fork of the Oniguruma (k.k, 2024) library, which is a Spencer-style backtracking regex engine.
ReDoS Defense: Until Ruby 3.2 (2022), Ruby’s port of Onigmo did not have built-in mitigations against ReDoS (issue 5820). In Ruby 3.2, the developers added two ReDoS defenses (rub, [n. d.]; Cha, [n. d.]). First, based on a community proposal, [^2] https://bugs.ruby-lang.org/issues/19104 Ruby implemented Davis et al.’s memoization technique for Spencer’s algorithm as a defense against ReDoS attacks (Davis et al., 2021). Interestingly, the Ruby developers used full memoization rather than selective memoization despite the higher space complexity. Ruby does not apply memoization when E-regex features are encountered, instead offering a timeout mechanism like C#. The memoization defense is on by default.
Engineering Effort: The relevant pull requests added a total of 1100 LOC, compared to the previous Ruby 3.1.6 engine’s footprint of 23384 LOC. The ratio is 5%.
4.1.3. C# (.NET)
Original Algorithm: The first two regex engines in .NET used backtracking algorithms (Bac, 2023; Reg, 2021). One was a simple opcode interpreter. The other used an optimized machine code representation like V8’s (Bac, 2023; Toub, 2023; Com, 2021; Gutierrez, 2004).
ReDoS Defense: Similar to Ruby, in .NET v4.5 (2012), .NET introduced a timeout mechanism to prevent excessive backtracking (Wha, 2023). The timeout is checked after at most $O(n)$ steps, where $n$ is the length of the input, balancing fidelity with overhead (Toub, 2022; Bes, 2023). In .NET 7 (2022), Microsoft added a non-backtracking regex engine (Toub, 2022). This engine is based on Moseley et al.  (Moseley et al., 2023) and Saarikivi et al. (Saarikivi et al., 2019), with its core algorithm using Brzozowski derivatives and guaranteeing linear-time matches. However, both the timeout mechanism and the non-backtracking engine are opt-in features.
Engineering Effort: The new engine comprises 5417 LOC. The existing .NET codebase for regexes was 15597 LOC (including multiple engines). Hence, the new engine expands the existing regex engine codebase by 35%. Microsoft researchers also created two research papers.
4.1.4. Java (OpenJDK)
Original Algorithm: OpenJDK Java’s previous regex engine implemented an NFA-based Spencer-style backtracking algorithm for regex matching (Pat, [n. d.]).
ReDoS Defense: As of Java 22 (2024), OpenJDK Java’s regex engine does not document defenses against ReDoS. Despite the absence of official documentation, in the OpenJDK source code repository and issue tracker, we identified that some performance updates related to ReDoS issues were introduced in Java 9 (2016). [^3] https://github.com/openjdk/jdk/commit/b45ea89 The OpenJDK maintainers introduced a bounded caching mechanism, improving problematic behavior for some ReDoS scenarios. This ReDoS defense is enabled by default and affects all OpenJDK Java versions $\geq$9.
Engineering Effort: Before the update, the regex-related codebase consisted of 4823 LOC. The memoization patch added 1712 lines (including empty lines and docstrings) to the Java 9 source code, which accounted for a 35% expansion.
4.1.5. Other ReDoS-Vulnerable Engines
PHP (Zend Engine), Perl (perl5), and Python (CPython) are also known to be using Spencer-style backtracking regex engines (Friedl, 2006; Davis et al., 2019), which are vulnerable to ReDoS attacks. According to Davis et al. (Davis et al., 2021), PHP (PHP Documentation Team, 2025) and Perl utilize counter-based caps to prevent catastrophic backtracking as an alternative to C# and Ruby’s time-based caps. Python has neither built-in engine optimizations against ReDoS vulnerabilities nor a native timeout-like mechanism for regex evaluations. We have not found any documented updates or defenses against ReDoS vulnerabilities for these languages since the analysis of Davis et al. (Davis et al., 2021) in 2021.
4.1.6. Other ReDoS-Safe Engines
Two major programming languages, Rust and Go, have regex engines that were developed with ReDoS safety in mind (Rus, 2024a; Gor, [n. d.]). These engines use an explicit automaton representation that is simulated with Thompson’s algorithm (Cox, 2007). They also do not support E-regex features. They, therefore, offer linear-time match guarantees in the length of the input and the automaton. Turoňová observed that the construction of the automaton might be slow or the memory footprint problematic if the regex uses counters (i.e., bounded quantifiers) (Turoňová et al., 2022).
Although not included in our analysis, independent third-party regex engines such as RE2 (Google) (Goo, 2024) and Hyperscan (Intel) (Wang et al., 2019; Wang, [n. d.]) are well-known for their ReDoS resilience. RE2, a Thompson-style engine, incorporates DFA state caching optimizations (Moseley et al., 2023). While generally ReDoS-safe and non-backtracking, RE2 can, in rare cases, exhibit backtracking behavior (Davis et al., 2021). Similarly, Hyperscan is another efficient, ReDoS-safe engine, utilizing hardware acceleration alongside Glushkov’s NFA construction (Glushkov, 1961). Neither engine supports E-regex features.
In conclusion, these developments in regex engines demonstrate significant progress in mitigating ReDoS risks but also highlight ongoing challenges in balancing the trade-off between safety and usability. Modern regex engines mitigate ReDoS with non-backtracking algorithms like memoization, Thompson-style simulation, or Brzozowski derivatives to assure linear-time matching. However, the benefit comes at the cost of reduced expressiveness. Non-backtracking engines lack support E-regex features. Defenses often supplement with timeouts or fallback mechanisms for E-regexes to maintain compatibility with existing codebases.
4.2. Updated Measure of ReDoS Vulnerabilities
As indicated in Table 5, several major regex engines have recently been updated to mitigate or eliminate ReDoS. We, therefore, wondered what proportion of ReDoS vulnerabilities have been addressed by state-of-the-art engines.
We answered this question through measurements on a sample from the Davis et al.’s (Davis et al., 2019) polyglot regex corpus, comprising $\sim$6K SL-regexes extracted from open-source repositories across many programming languages. We filtered this corpus for all regexes predicted to be SL by vuln-regex-detector (dav, 2024), a tool developed by Davis et al. (Davis et al., 2018a), which relies on (Rathnayake and Thielecke, 2017; Wüstholz et al., 2017; Weideman et al., 2016; Shen et al., 2018). We evaluated a sample of 500 exponential-time regex candidates (out of 3K examples) and 500 polynomial-time regex candidates (out of $\sim$100K examples). For these regexes, we generated attack inputs using, again, the vuln-regex-detector tool.
Then, we evaluated each regex-input pair on the regex engines described in Table 5, comparing the performance of the engine version before any changes to that of the most recent version. If ReDoS defenses were available but not enabled by default in the latest version—as in the case of JavaScript and C#—we activated them. Following a parallel methodology that has been used in previous work (Davis et al., 2019, 2018a), we define a regex’s behavior as:
  • Exponential: The processing time exceeds five seconds with fewer than 500 pumps of an attack string.
  • High-Polynomial: The processing time exceeds five seconds with 500 or more pumps of an attack string.
  • Low-Polynomial: The processing time exceeds 0.5 seconds on an attack string but does not meet the exponential or high-polynomial behavior criteria.
  • Linear: The regex match never times out on the attack string.
Figure 5 shows our results. For each engine, we depict the number of regexes that exhibited SL behavior (as a fraction of the predicted input) in the pre- and post-ReDoS mitigation versions. Not all regexes were supported in all engines, so the candidate pool size is indicated for each engine. Observe that for the engines with mitigations, there is a substantial decrease in the number of viable SL-regexes—in other words, these mitigations appear to be effective in addressing the common causes of SL behavior. Notably, exponential behavior is resolved in C# and JavaScript but persists in Ruby and Java, while polynomial behavior continues to manifest in JavaScript and Java.
Figure 5. Performance of regex matching in old vs. latest engines in different programming languages.
5. Discussion
In this section, we distill and contextualize the main findings of our study, integrating insights from our literature review (Section 3) and engineering review (Section 4). For each idea, we discuss concrete future work directions for the community to explore.
(1) ReDoS is a peculiar vulnerability. Different from more traditional vulnerabilities, e.g., SQL injections, ReDoS stands out as a contextual issue, more similar to micro-architectural attacks, where exploitability and impact are heavily dependent on the underlying system’s architecture and deployment conditions. For example, a regex pattern vulnerable to catastrophic backtracking in Node.js v15 is rendered safe in v20 due to advancements in regex engines (Section 4). When detecting ReDoS vulnerabilities, it is not enough to analyze the code of a web application, without considering its deployment configurations. For future static analysis work on ReDoS, e.g., that targets multiple languages (Hassan et al., 2023), we recommend that they consider and model the language-, engine-, and platform-specific details and differences with respect to ReDoS. For example, a super-linear regular expression might be more serious in a single-threaded web application architecture like Node.js than in inherently parallel ones like Go. For dynamic and hybrid analyses, we recommend a system perspective, similar to Shan et al. (Shan et al., 2017), in which the code under test is deployed under realistic conditions, where potential defenses like web application firewalls are also considered.
(2) Strong assumptions about attackers. We find that prior work often makes strong assumptions about attackers (Section 3.2), e.g., they can inject arbitrarily large payloads, or they can control the input to any regular expression. Also, many papers neglect to deploy defenses or rely on simple rules of thumb such as that a regular expression match should not take longer than one second (Section 3.3). We think that future work should challenge these assumptions by performing empirical studies with real systems, e.g., to analyze how many regular expressions can be reached by attacker-controlled inputs. We also need more empirical studies to help the community understand how modern applications are deployed in the wild, e.g., in which specific cloud configurations and what defenses they use. Finally, future ReDoS research should perform realistic evaluations by measuring the broader impact of attacks, including system slowdowns and user experience. To aid with this, in Appendix C, we propose wrk-redos, a tool for simulating realistic ReDoS scenarios by limiting attackers’ capabilities such as the request’s length and attacks’ bandwidth. By evaluating vulnerabilities under practical conditions, tools like wrk-redos can help developers understand risks more effectively and implement appropriate mitigations.
(3) Developers’ push back. To the best of our knowledge, there has been no public description of a denial-of-service attack that exploited ReDoS, so the risk posed by these vulnerabilities is still unclear, beyond academic measurement studies (Staicu and Pradel, 2018; Barlas et al., 2022). Adding to the confusion, only a small fraction of the papers we analyzed (Section 3) discuss such attacks, often in unrealistic setups. We believe that these unclear risks together with problematic security assumptions are the culprits for ReDoS vulnerabilities being a contentious topic among developers. Snyk has stated that ReDoS is the most neglected vulnerability type (ope, [n. d.]), implying that developers often ignore such findings. Also, engineers at GitHub consider it a low severity vulnerability, “not particularly serious, but easy to create by accident, obscure to understand, and sometimes tricky to fix” (Backhouse, 2023). More strident voices argue that ReDoS vulnerabilities are “malicious noise” and just lead to security fatigue (Woodruff, 2022). Several ReDoS CVE advisories have been disputed by the maintainers (e.g., CVE-2023-39663, CVE-2022-42969, CVE-2019-11391), indicating disagreement between maintainers and reporters, also reflected by negative comments in GitHub issues (Section D.4).
The blame for this perspective may lie with the cybersecurity academic community. As noted in Section 3, researchers have performed large-scale measurements of ReDoS as well as produced new and more effective detectors for vulnerable regexes. Their measurement instruments have been integrated into popular security scanners, e.g., the integration of safe-regex into eslint. However, the resulting scanners analyze only the regex, without considering reachability constraints as well as other input constraints. When academics apply these tools, they (should) filter false positives based on the regex’s context. When practitioners do so, they may make much ado about nothing. This suggests that empirical studies of academic tools uptake is of interest, as well as prompting the research community to reflect on direct access to research tooling and more developer-friendly releases. More importantly, future work in the field should perform user studies with developers to understand their pain points with respect to ReDoS (tools), security assumptions, and deployment configurations. In this way, we can better align the community’s assumptions with those of practitioners.
(4) Engines have evolved significantly. It is encouraging that modern language runtimes changed their implementation in response to prior work on ReDoS (Section 4). These changes were often driven or informed by researchers, e.g., Ruby’s adoption of Davis’s proposals (Davis, 2019; Davis et al., 2021). [^4] See https://bugs.ruby-lang.org/issues/19104. We believe that this is a success story like few in our community that shows how researchers and practitioners can collaborate effectively. Our study shows that modern regex engines have evolved significantly, incorporating advanced techniques such as Thompson’s algorithm, memoization, and resource capping to mitigate ReDoS. However, these solutions are often complex and rarely as clean as their theoretical counterparts. Additionally, there are often redundant implementations that are triggered on demand by the engine, when matching a regular expression. Hence, future work should investigate the deployed fixes to find potential bugs or misbehaving corner cases. We advocate for comparative studies like ours Section 4, in which multiple engines are investigated on the same regex/dataset to study the advantages and disadvantages of the deployed defenses. The research community should also propose automatic solutions for migrating to patched engines. More importantly, research on ReDoS should consider these advancements to ensure findings remain relevant to contemporary systems.
(5) Performance bugs vs. security vulnerabilities. Many of our findings also apply to other algorithmic complexity vulnerabilities, e.g., Zip bombs. We note that while the possibility of weaponizing slow computation is real, attackers still seem to prefer brute-force, volume-based DDoS attacks. Researchers should aim to understand why that is the case and, simultaneously, continue to develop defenses against such potential future attacks. We also believe that there are many insufficiently-explored, cross-disciplinary opportunities for ideas that aim to bridge the gap between performance, measurement, and security communities. For example, we argue that most performance bugs, e.g., redundant loop traversals (Olivo et al., 2015b), can probably be lifted to a DoS attack when malicious intent is assumed. Thus, future work, should investigate in more detail the relation between performance bugs and CPU-based DoS attacks, in particular, in modern web applications.
6. Conclusion
In this work, we perform a multi-faceted study of the recent academic work on ReDoS and the new engineering realities in the mainstream programming languages, after initial fixes were deployed in most of them. We report on concrete examples of language runtimes that adopted defenses, which appear to be motivated by academic research. When studying the efficacy of these defenses, we find that they mitigate most ReDoS payloads, but they still leave significant room for attackers to maneuver. By surveying the academic work on ReDoS, we find that many papers in this domain use weak definitions and strong threat models, often considering any slowdown caused by regular expression matching as a security vulnerability. We advocate for future work to consider the recently-deployed fixes in mainstream programming languages and to evaluate ReDoS attacks under realistic deployment conditions, e.g., against a web application deployed in the cloud using reversed proxies and redundancy. We also propose restricting the attackers’ capabilities to realistic payload sizes and modest bandwidths that comply with the asymmetric DoS setting. More importantly, future work should study the more general problem of weaponizing slow computation, which the attackers might leverage in the near future.
7. Ethics Considerations
In our judgment, this work presents an acceptable ethical risk-reward tradeoff. Specifically:
  • The literature review (Section 3) and engineering review (Section 4) describe and systematize only publicly-available resources.
  • Any measurement tool such as wrk-redos (Section C.1) poses dual use dilemmas: it helps improve systems but may also aid attackers. We believe its benefits outweigh the risks.
  • In our evaluation of wrk-redos (Section C.2), we only rely on known/patched vulnerabilities, and we only simulate ReDoS attacks against web servers that we control.
8. Compliance with Open Science Policy
An artifact accompanies this paper at https://anonymous.4open.science/r/wrk-DoS-BD74. The artifact includes (1) the dataset of super-linear regexes with their corresponding attack strings, and (2) the source code and usage manual for the realistic ReDoS attack simulation framework proposed in Appendix C.
Acknowledgements.
This work was supported in part by the Sponsor National Science Foundation (NSF) under grant #Grant #2135156.
Appendix A Regex Feature Reference
Table 6 provides a comprehensive reference of regex features and notation, summarized from the PCRE2 specification (Hazel, 1997). The table categorizes regex features into two groups: K-regexes and E-regexes. E-regexes are also subcategorized. Each feature is associated with its syntactic notation.
Table 6. Full set of features and notation of regexes, compiled from (Hazel, 1997). $R$ is a sub-pattern.
Abbrev. Feature Notation
K-regexes
CAT X followed by Y $R_{1}R_{2}$
KLE Zero-or-more rep. $R*$
OR Logical OR $R_{1}|R_{2}$
E-regexes
Capture Groups
CG Capture group $(R)$
NCG Non-capture group (?$:R)$
[…]
Quantifiers
ADD One-or-more rep. $R+$
QST Zero-or-one rep. $R?$
DBB Range repetition $R\{m,n\}$
LWB At-least-$m$ rep. $R\{m,\}$
SNG Exactly-$n$ rep. $R\{n\}$
Character Classes
CCC Custom char. class $[aeiou]$
RNG Char. range $[a-z]$
NCCC Negated class [^$[\textasciicircum{aeiou}]$
ANY Any char. .
WSP Whitespace $\setminus{}s$
DEC Numeric $\setminus{}d$
WRD Word $\setminus{}w$
NWSP Non-whitespace $\setminus{}S$
NDEC Non-numeric $\setminus{}D$
NWRD Non-word $\setminus{}W$
VWSP Vertical space $\setminus{}v$
Zero-width Assertions
STR Start-of-string/line ^$\textasciicircum R$, $\setminus{}\text{A}R$
END End-of-string/line $R\$$, $R\setminus{}\text{Z}$
[…]
PLA Pos. lookahead (?$=R)$
NLA Neg. lookahead (?$!R)$
PLB Pos. lookbehind (?<${\textless}=R)$
NLB Neg. lookbehind (?<${\textless}!R)$
Backreferences
BKR Numeric backref. $(R)\ldots\setminus{}1$
[…]
Backtracking Controls
LZY Non-greedy rep. $R*?,R+?,R\{m,n\}?$
ATM Atomic group (?>${\textgreater}R)$
POS Possessive quantifier $R++,R*+$
Appendix B ReDoS in Practice
This section provides a detailed account of the study described in Section 2.3.2. It outlines the methodology employed to investigate ReDoS vulnerabilities, including the criteria for selecting relevant data and the dimensions of analysis. Additionally, the section presents the key findings and results of the study
B.1. Methodology
B.1.1. ReDoS in the National Vulnerability Database
We performed a comprehensive analysis of NVD data from 2014 to 2023, focusing on CVEs associated with OWASP Top-10 and ReDoS-related vulnerabilities, using the classifications available on the OWASP website [^5] https://owasp.org/www-project-top-ten/. To ensure that ReDoS CVEs were accurately identified, we employed a systematic two-stage keyword search:
  • Stage one keywords: regex, regular expression, regexp
  • Stage two keywords: algorithm, backtrack, uncontrolled, repetition, repeat, infinite, denial of service, dos, infinite loop, algorithmic complexity
This method effectively identified mislabeled ReDoS CVEs, with our checks confirming no false positives in the sampled data. We further categorized the CVEs by their respective ecosystems, using specific keywords for each:
  • npm: npm, node.js, nodejs, javascript, js
  • PyPI: pypi, python, pip, django, flask
  • Gems: rubygems, ruby, gems, rails
  • Maven: maven, java, spring
  • Packagist: packagist, composer, php, laravel
  • Cargo: cargo, rust, crates.io, rust-lang
Our verification of a sample from each ecosystem showed no false positives, confirming the accuracy of our categorization.
B.2. Results
B.2.1. ReDoS in the National Vulnerability Database
Our analysis of CVE data from the NVD over the past 10 years has provided insights into the incidence of OWASP Top-10 and ReDoS-related CVEs. These trends are illustrated in Figure 6, with a specific focus on ReDoS-related CVEs presented separately in Figure 7. Additionally, we explored the distribution of OWASP Top-10 and ReDoS CWEs across different ecosystems, as shown in Figure 3. This visualization also includes the average occurrence of each CWE type throughout all ecosystems.
Figure 6. Number of OWASP Top-10 and ReDoS-related CVEs (2014-2023).
Figure 7. Number of ReDoS-related CVEs (2014-2023).
Appendix C wrk-redos Framework
The wrk-redos framework is designed to provide a practical, simulation-based approach to evaluate the severity of ReDoS vulnerabilities. Although many existing studies analyze ReDoS from a theoretical perspective, real-world scenarios often involve complex interactions between the regular expression engine, the application, and the underlying infrastructure. wrk-redos addresses this gap by offering tools and methodologies to simulate realistic attack scenarios, allowing developers and researchers to understand the impact of ReDoS vulnerabilities on system performance.
This framework is beneficial for the following purposes:
  • Benchmarking quality of service (QoS) degradation during a ReDoS attack.
  • Validating the severity of vulnerabilities found in regexes, especially in scenarios with third-party dependencies.
  • Providing actionable insights to mitigate the risks associated with vulnerable regex patterns.
The following subsections outline the use cases, threat model, and implementation details of the framework (Appendix C.1), along with the results of simulations (Appendix C.2) and an analysis of their applicability in realistic environments (Appendix C.3).
C.1. Realistic Simulation of ReDoS Attacks
As shown in Section 3, academic work on ReDoS can benefit from more realistic threat models and availability-oriented evaluations, to better communicate the ReDoS-associated risks. To this end, we present a framework for assessing the severity of vulnerable regular expressions via attack simulations. We first define the main considered use cases, the threat model and the ReDoS definition, then present wrk-redos, our prototype for performing measurements in the proposed framework.
Use Cases. As discussed in the introduction, developers are often overwhelmed by the number of supposedly-vulnerable regular expressions in their code base, in particular, in code they did not write, e.g., shipped with third-party dependencies. To confirm the severity of such security findings, we propose simulating actual ReDoS attacks and benchmark the corresponding degradation in quality of service, as perceived by concurrent users. In this way, developers can understand the attacker’s leverage in an asymmetric DoS, enabled by the problematic regular expression.
Considering the current state of regex engines discussed in Section 4, we note a second usage scenario for our framework. It can assist developers with understanding the effect of migrating their application to a different engine version, from a ReDoS resilience perspective. Concretely, developers can perform attack simulations before and after changing the runtime version and measure how the attacker’s advantage changes as a result of this migration. Since regex engine migration carries risks, notably that of semantic changes, wrk-redos can help engineers assess the risk-reward tradeoff here.
Threat Model. We assume that attackers and users can interact simultaneously with the target system via two endpoints: one that is benign and one that is vulnerable to ReDoS attacks. Attackers have limited bandwidth and know how to create ReDoS payloads, but we do not assume any special deployment features like redundancy or intrusion detection systems. This model restricts attackers from sending excessively large payloads and does not presume a specific regex engine type, which aligns more closely with realistic scenarios. Under this threat model, we propose an alternative definition for ReDoS vulnerabilities that we believe is more in line with the practitioners’ expectations:
Definition 0.
A problematic regular expression is considered a vulnerability if and only it is accompanied by an attack simulation with the following characteristics:
  • C1: The vulnerability is exercised in a realistically-deployed web application, e.g., in the cloud.
  • C2: The attackers’ bandwidth is restricted to a reasonably low value, in line with the asymmetric setup.
  • C3: The simulation shows that concurrent benign users experience dropped requests or extremely high latency.
This definition has certain advantages over previous ones proposed in the literature. It discards minor slowdowns caused by ReDoS issues, which cannot be leveraged for an asymmetric attack. It emphasizes the server’s QoS degradation, as perceived by benign users, instead of merely measuring the slowdown produced by individual payloads. Moreover, it requires critical availability consequences like dropped requests for an issue to be considered a vulnerability. We believe that this will help alleviate the security fatigue regarding ReDoS and prioritize harmful ReDoS issues. Most importantly, our definition mandates that the web application needs to be deployed in a setup that closely resembles a real production environment. In this way, certain deployment-time components such as the reversed proxy might absorb some of the malicious ReDoS traffic and inadvertently prevent the attack.
Figure 8. Overview of wrk-redos, an attack simulation framework for a realistic assessment of ReDoS findings.
Simulation Framework. Considering the criteria we established for a successful attack, we propose wrk-redos a simulation tool that measures the effectiveness of ReDoS attacks leveraging a target regular expression. Figure 8 depicts how wrk-redos operates. wrk-redos takes the user’s simulation configuration, such as attack bandwidth and window, measurement bandwidth, and experiment duration. It then launches two concurrent workloads: (i) a benign one that is triggered first and continuously benchmarks the target server, i.e., sends high-bandwidth measurement requests, (ii) a malicious one that sends ReDoS payloads in a restricted time window, at a predefined rate. wrk-redos communicates with these processes, manages the data configuration, and gathers output metrics, such as throughput, latency, and errors for benign users. Adopting a remote benchmarking approach, wrk-redos highlights the attack’s impact on benign users (C3), measuring the system’s capability to handle both legitimate and malicious traffic concurrently. Nonetheless, to comply with the other two simulation characteristics in the definition above, users of wrk-redos must ensure that they configure attacks with sufficiently low bandwidth (C2) and that the target application is deployed under realistic conditions (C1). Similar to related work such as Rampart (Meng et al., 2018), Shan et al. (Shan et al., 2017), and Olivo et al. (Olivo et al., 2015a), our framework performs short attack bursts and observes the server’s behavior before and after the attack. This method enables asymmetric attacks, similar to low-rate denial of service and Shrew attacks(Zhijun et al., 2020), both known to bypass existing DoS defenses. Moreover, wrk-redos utilizes wrk2 (ten, 2024), a popular HTTP benchmarking tool (Fuerst et al., 2020; Unnikrishnan, 2023; Ding et al., 2023; Gouveia et al., 2020; Jacquet et al., 2023; Maissen et al., 2020) to simulate both benign and malicious traffic. This allows us to maximize the amount of traffic sent from one machine and reuse a mature measurement infrastructure adopted both by practitioners and academics. wrk2 follows standards set in benchmarking literature to ensure accurate, stable, and relevant measurements.
C.2. Experiments
Table 7. Performance metrics for 10 NPM and PyPI libraries at input sizes 25, 1000, and 5000. Metrics include 99.99th percentile latency ($p_{99.99}$), mean ($\bar{x}$), timeout requests ($N_{u--sw}$), and failed requests ($N_{failed}$).
Library Input size = 25 Input size = 1000 Input size = 5000 Input size = 5000 (RE2) $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $N_{u--sw}$ $N_{failed}$ black 561.49 128.66 0 0 178.27 132.70 3401 (75.5%) 50 777.65 136.52 3251 (72.2%) 50 247 (13.7%) 0 celo 173.76 124.52 3401 (75.5%) 50 412.93 111.14 3401 (75.5%) 50 151.77 128.00 3301 (73.3%) 50 0 0 django-truncate 438.82 113.41 0 0 479.60 154.64 0 0 12801.15 2615.76 0 0 N/A N/A django-intcomma 426.59 111.62 0 0 189.72 122.68 0 0 7332.25 1121.46 0 0 N/A N/A semgrep 9644.78 1675.03 0 0 459.06 120.57 3301 (73.3%) 50 543.81 139.27 3396 (75.4%) 50 0 0 charset 574.81 129.82 0 0 483.45 129.83 0 0 5496.47 134.67 0 0 0 0 node-email-check 342.45 137.12 3301 (73.3%) 50 487.60 133.98 0 0 621.69 117.48 0 0 0 0 cookiejar 704.84 122.32 0 0 185.87 118.65 0 0 5238.85 847.67 0 0 0 0 lodash 437.11 112.79 0 0 539.86 129.12 0 0 197.13 126.80 0 0 0 0 semver 652.86 110.85 0 0 1407.34 164.78 0 0 10154.02 1797.97 0 0 0 0
Table 8. Overview of npm and PyPI libraries used in the study, including their versions, associated CVEs, programming languages, and severity scores.
Library Version CVE Language Severity
black 24.1.0 CVE-2024-21503 Python 5.3
celo 1.0.0a5 CVE-2022-42966 Python 5.9
django-truncate 3.2.0 CVE-2023-43665 Python 5.9
django-intcomma 3.2.0 CVE-2024-24680 Python 5.9
semgrep 1.23.0 CVE-2023-32758 Python 7.5
charset 1.0.0 CVE-2017-16098 Nodejs 7.5
node-email-check 1.0.0 CVE-2023-39619 Nodejs 7.5
cookiejar 2.1.3 CVE-2022-25901 Nodejs 5.3
lodash 4.17.16 CVE-2020-28500 Nodejs 5.3
semver 5.1.1 CVE-2022-25883 Nodejs 5.3
To illustrate the utility of wrk-redos, we analyzed previously reported ReDoS vulnerabilities in open-source libraries by conducting equivalent attacks of varying intensities. We selected ten libraries from the npm and PyPI ecosystems, all with reviewed vulnerability reports in the GitHub Advisory Database. These libraries were chosen based on their severity scores and available proof-of-concept payloads. Details about the tested libraries and their versions are provided in Table 8. Despite their popularity—eight libraries with over 850,000 weekly downloads and seven with more than 3.5 million—these libraries are still susceptible to security issues. We used both Heroku’s basic dyno (512 MB memory, 1 CPU) and an AWS EC2 instance (1 GB memory, 1 CPU) to set up our sample applications. The EC2 instance ran custom Docker applications to implement the second use case described in Section C.1, which involved replacing Python’s re module with RE2 and Node.js’s regex engine with RE2, instead of the default V8 engine. However, we could not perform the same measurements for django-intcomma due to a stack overflow caused by RE2’s internals. This setup enabled us to simulate and measure the impact of migrating to a ReDoS-safe engine, comparing ReDoS resilience and performance before and after the engine swap. We created Express and Flask applications to process user input through these vulnerable libraries, with dedicated API endpoints for each. Malicious payloads were sent via headers to test the servers’ vulnerability to attacks.
Figure 9. Throughput and latency for benign users for django-truncate library at payload sizes 25, 1000, 5000, and 10000.
Table 7 presents the results of our simulation experiments using three payload sizes (25, 1000, and 5000 characters), with an attack burst lasting 10 seconds at a rate of 100 req/s, and a measurement window of 60 seconds at the same rate. We calculate and list various metrics that demonstrate the resilience of the target systems under these attacks. As the intensity of the attack increases, we observe a consistent rise in latency metrics (average and tail). Additionally, the last two columns of the table show the number of timeouts and failed requests with the 5000-character payload when RE2 was enabled in both Node.js—V8 and Python. Despite charset having a high severity score of 7.5, its impact on performance metrics such as average and median latencies was minimal, with increases only in the millisecond range except for some outliers at the $99.99th$ percentile ($p_{99.99}$). Conversely, semver, with a lower severity score of 5.3, seriously reduced server throughput to less than 10% with an input size of 5000 characters and significantly increased the average latency of benign requests for 1000 characters. These observations suggest that the severity scores provided in advisory databases might not consistently reflect their effect on a system’s vulnerability to DoS attacks. Additionally, different advisory databases may assign varying severity scores to the same vulnerability, as seen with node-email-check, where GitHub Advisory Database rates it as a high vulnerability (7.5), whereas Snyk.io considers it medium severity (5.3). Interestingly, celo exhibited a complete server shutdown with an input length of only 25 characters but showed no adverse effects with larger inputs of 1000 and 5000 characters. This resilience at larger sizes can be attributed to the input restriction implemented in the celo library, which sanitizes input by ensuring the email length is greater than 5 and less than 61 characters, as shown in Figure 10. While this approach effectively filters out malicious inputs longer than 61 characters and prevents large-payload attacks, it does not prevent Regular Expression Denial of Service (ReDoS) attacks with smaller, targeted inputs exploiting the vulnerable expression.
1function validateEmail (email) {
2 if (typeof email === ’string’ && email.length > 5 && email.length < 61 && validateRegex.test(email)) {
3 return email.toLowerCase();
4 } else {
5 return false;
6 }
7}
Figure 10. Email validation code snippet from Celo library
Let us now consider the case of django-truncate in more detail. Figure 9 illustrates the throughput and latency of the Django application under various attack scenarios. In Django version 3.2, the method django.utils.text.Truncator is vulnerable to ReDoS, when used with html=True. In our setup, malicious data is sent by including it in an HTTP header — a common method for data transmission to the server — to exploit this vulnerability. The graph depicts the server’s response to different simulated attack scenarios with payloads of lengths 25, 1000, 5000, and 10000. We observe an increase in both average and median latency as the payload sizes increase from 25 to 1000, indicating a degradation in server performance during the attack. Moreover, with a payload of 5000, there is a significant reduction in server throughput, approximately to 10%, which signifies a substantial impact of the attack. However, this throughput rebounds to baseline levels shortly after the cessation of the attack. Conversely, the 10000 payload scenario deviates from the expected outcome, exhibiting no discernible effect on server performance. This anomaly arises because the payload size exceeds Heroku’s header limit, resulting in a rejection of the malicious requests by the server’s infrastructure. Thus, despite the attacker dispatching larger payloads, the server remains unaffected, due to an intrinsic safeguard against excessively large payloads within the platform. This supports our claim that ReDoS findings should be evaluated in realistic scenarios, where such complexities are considered.
C.3. Analysis of Simulation Results
When simulating attacks like ReDoS, it is essential to employ a multi-faceted measurement strategy that goes beyond single metrics such as tail latency. This avoids the over-simplification trap of a single viewpoint, as discussed below.
Detailed Analysis of Metrics. Let us consider the case of semgrep in Table 7 to illustrate how relying solely on tail latency can sometimes paint an incomplete picture of an attack’s true impact. With input sizes of 25 versus 1000, the data suggests a disproportionate effect on server performance at smaller input sizes. Initially, it appears more disruptive due to higher latency spikes. Yet, this view does not investigate the server’s ability to recover quickly once the attack ceases, which is crucial for assessing the system’s resilience.
As shown in Figure 11, the effect of a ReDoS attack on the semgrep library using input sizes of 25 and 1000 characters highlights this issue. For an input size of 25, the throughput decreases to almost 10%, but it recovers immediately once the attack ceases. However, with an input size of 1000, the server fails to recover within the experiment window, demonstrating a significant vulnerability to larger ReDoS attacks. Even though the reported tail latency is lower due to incomplete data, the actual impact is significantly more severe since the server remains non-operational.
Figure 11. Server throughput over time for input sizes of 25 and 1000 using the Semgrep library.
Figure 12. Throughput and latency for benign users at 500 req/s bandwidth. The increasing latency and fluctuating throughput indicate bandwidth exceeding server capacity, leading to an inadvertent DDoS attack.
Effects of Payload Size. We note that there is often a key threshold beyond which increasing the payload size has no additional negative impact. For instance, comparing input sizes of 1000 and 5000 for semgrep in Table 7 shows negligible differences in the observed metrics. This indicates that once a payload is sufficiently large to overwhelm the server, further increases do not exacerbate the attack’s effect.
Setting Appropriate Bandwidth. When evaluating the severity of a ReDoS exploit, it is crucial to ensure that the simulation itself does not inadvertently launch a traditional DoS attack on the server. This can be avoided by properly setting the attack bandwidth in the benchmarking tool; otherwise, the results fail to show the effect of the ReDoS payload, but of the high-bandwidth traffic. Figure 12 illustrates this issue, showing the consequences of running wrk-redos at an attack rate of 500 requests per second, which exceeds the server’s capacity. This is indicated by the fluctuations in throughput and the increase in latency even before the actual attack begins. As a result, the observed performance degradation can be attributed mainly to the high traffic volume, instead of the ReDoS payloads. Therefore, it is essential to calibrate both the benign and attack bandwidths to accurately assess the impact of a ReDoS attack.
Stable Measurement Conditions. Preparing and calibrating the system before conducting attack simulations is essential. This involves stabilizing both throughput and latency with benign loads to ensure that the baseline performance is reliable and reflective of typical operational conditions. wrk-redos facilitates this by simulating benign conditions, at first, to verify the system’s stability before launching the attack simulation, helping to ensure that the subsequent measurements accurately reflect the attack’s impact.
Appendix D Other Tables and Figures
Table 9. Regex engine implementations selected for evaluating ReDoS defenses
PL Selected Impl. of the PL [Source Code] Other Alternatives Reason for Selection
JavaScript Node.js—V8 (Nod, 2024; V8V, 2024) Deno, Bun, SpiderMonkey, JavaScriptCore Dominates server-side JavaScript usage (98.8% market share) (Dis, [n. d.])
Ruby MRI/CRuby (Rub, 2024) JRuby, Rubinius, mruby, RubyMotion Reference implementation (Rub, [n. d.])
C# .NET (Dot, 2024) Mono Reference implementation (Lan, 2024)
Java OpenJDK (Ope, 2024) Amazon Corretto, Azul Zulu, Liberica JDK, Eclipse Adoptium… Most widely used JDK of 2024 with a 20.8% usage rate (202, [n. d.])
PHP Zend Engine (Php, 2024) HHVM, PeachPie, Quercus, Parrot Standard PHP interpreter is driven by the Zend Engine (PHP, [n. d.]; php, [n. d.])
Perl perl5 (Per, 2024) N/A (for Perl 5) Only implementation for Perl 5 (Per, [n. d.])
Rust rustc (Rus, 2024b) mrustc, gccrs Official and only fully functional compiler for Rust (rus, [n. d.])
Go gc (Gol, 2024) gccgo, gofrontend, TinyGo, GopherJS, yaegi Official compiler included in Go releases (Int, [n. d.])
Python CPython (Pyt, 2024) PyPy, Stackless Python, MicroPython, CircuitPython, IronPython, Jython Reference implementation (Pyt, [n. d.])
PL: Programming Lanuage, Impl.: Implementation (runtime environment, engine, interpreter, or compiler)
Table 10. List of all works included in this SoK, sorted by year of publication.
Paper Year Venue Kirrage et al. (Kirrage et al., 2013) 2013 NSS Weideman et al. (Weideman et al., 2016) 2016 CIAA Davis et al. (Davis et al., 2017) 2017 EuroSec Rathnayake et al. (Rathnayake and Thielecke, 2017) 2017 arXiv Wüstholz et al. (Wüstholz et al., 2017) 2017 TACAS Petsios et al. (Petsios et al., 2017) 2017 CCS Staicu et al. (Staicu and Pradel, 2018) 2018 USENIX Security Davis et al. (Davis et al., 2018a) 2018 ESEC/FSE Shen et al. (Shen et al., 2018) 2018 ASE Noller et al. (Noller et al., 2018) 2018 ISSTA Davis et al. (Davis et al., 2018b) 2018 USENIX Security Davis et al. (Davis et al., 2019) 2019 ESEC/FSE Demoulin et al. (Demoulin et al., 2019) 2019 USENIX ATC Li et al. (Li et al., 2020) 2020 ASE Turoňová et al. (Turoňová et al., 2020) 2020 OOPSLA Liu et al. (Liu et al., 2021) 2021 IEEE S&P Li et al. (Li et al., 2021) 2021 USENIX Security Bai et al. (Bai et al., 2021) 2021 IEEE S&P Davis et al. (Davis et al., 2021) 2021 IEEE S&P Turoňová et al. (Turoňová et al., 2022) 2022 USENIX Security Barlas et al. (Barlas et al., 2022) 2022 ICSE Kluban et al. (Kluban et al., 2022) 2022 AsiaCCS McLaughlin et al. (McLaughlin et al., 2022) 2022 USENIX Security Chida et al. (Chida and Terauchi, 2022) 2022 IEEE S&P Li et al. (Li et al., 2022) 2022 USENIX Security Bhuiyan et al. (Bhuiyan et al., 2023) 2023 ICSE Wang et al. (Wang et al., 2023) 2023 IEEE S&P Parolini et al. (Parolini and Miné, 2023) 2023 Science of Computer Prog. Hassan et al. (Hassan et al., 2023) 2023 IEEE S&P Tandon et al. (Tandon et al., 2023) 2023 RAID Su et al. (Su et al., 2024) 2024 USENIX Security Siddiq et al. (Siddiq et al., 2024a) 2024 ICPC Siddiq et al. (Siddiq et al., 2024b) 2024 ICSE Kluban et al. (Kluban et al., 2024) 2024 TOPS
D.1. Regex Engine Selection Summary
Table 9 provides a comprehensive overview of programming language, hence the regex engine implementations selected for our ReDoS defenses review (Section 4). For each programming language considered in our study, we list the chosen implementation, its alternatives, and the rationale behind our selection. In summary, this table reflects our effort to analyze the most relevant and widely-used regex engines in production environments, focusing on reference implementations or those with significant usage rate compared to alternatives where possible.
D.2. Included Works in the Systematization
Table 10 presents all the works included in the literature review in Section 3 of this SoK. The papers are sorted by year of publication and also include the publishing venue or conference. These works were selected based on the inclusion criteria described in Section 3, ensuring relevance to ReDoS-specific research. The selected papers represent the state-of-the-art in the field, covering key aspects such as detection, prevention, and mitigation strategies, and reflecting significant contributions to advancing the understanding of ReDoS.
D.3. Excluded Works in the Literature Review
Some closely related papers were excluded from the literature review because they do not focus on ReDoS-specific vulnerabilities or their evaluation lacks any ReDoS-specific aspects. These exclusions were necessary to ensure that the review remains precise and relevant to the study of ReDoS. For instance, (Cox, 2007; Varatalu et al., 2025; Turoňová et al., 2022; Shan et al., 2017) were excluded because they focused on broader denial-of-service (DoS) vulnerabilities or algorithmic complexity attacks. Other papers were excluded for evaluating performance issues in regex engines without demonstrating their impact on ReDoS vulnerabilities.
Table 11 highlights examples of these excluded works along with the reasons for their exclusion, showcasing the deliberate scope refinement undertaken to focus on ReDoS-specific contributions.
Table 11. Example works excluded with explanation
Paper Reason for Exclusion
Cox, R. (Cox, 2007) Evaluation does not address ReDoS-specific scenarios.
Varatalu et al. (Varatalu et al., 2025) Evaluation does not address ReDoS-specific scenarios.
Chattopadhyay et al. (Chattopadhyay et al., 2025) Evaluation does not address ReDoS-specific scenarios.
Turoňová et al. (Turoňová et al., 2022) Evaluation does not address ReDoS-specific scenarios.
Shan et al. (Shan et al., 2017) Addresses general denial-of-service (DoS) attacks.
Wei et al. (Meng et al., 2018) Addresses general denial-of-service (DoS) attacks.
D.4. Analyzing Developer Discussions on GitHub ReDoS Issues
To analyze discussions around ReDoS vulnerabilities, we utilized the GitHub REST API to search for issues containing the term redos in their title, body, or comments. This approach enabled the identification of both open and closed issues across various repositories, capturing a wide range of cases, including those actively being discussed and those already resolved. To focus on meaningful engagement, we filtered the results to include only issues with at least two comments, ensuring the dataset represented discussions with substantive input. After retrieval, we manually reviewed the comments within these issues to identify and extract relevant discussions about ReDoS vulnerabilities.
Table 12 showcases 10 examples of comments from selected issues, providing URLs for further exploration. These discussions often reveal that developers do not perceive ReDoS vulnerabilities as relevant or critical, leading to delays or reluctance in addressing the issues. While some developers prioritize security, others downplay the severity of these vulnerabilities, resulting in limited attention or resolution efforts. These trends underscore the challenges faced in addressing ReDoS issues in open-source projects.
Table 12. Selected ReDoS-Related Comments and URLs
Comment URL
D3 is almost never used to handle user input, so these ReDoS vulnerabilities are generally a huge waste of time. https://github.com/d3/d3/issues/3939
indeed, but since you’d have to be attacking yourself to trigger the ReDOS in npm, it’s not actually a vulnerability here. https://github.com/npm/cli/issues/7902
None of these seem like they could be exploitable in real world scenarios. Please let us know if you disagree. https://github.com/facebook/react-native/issues/46996
The few times there was an actual vulnerability, it was reported separately, and we released patches as soon as it was possible. You can always report real vulnerabilities here, but please do this if you understand the difference between a real vulnerability and a false positive. For example, a ”Regex DDOS attack” can never be a real vulnerability for a development-time tool. If you’re not sure, you’re welcome to ask in this thread, but please keep it brief and to the point so that the thread doesn’t become unreadable. https://github.com/facebook/create-react-app/issues/11174
This kind of low-quality finding was assigned a CVE ID without any significant cross-checking. https://github.com/pytest-dev/py/issues/287
So, a DoS can be caused by a package that’s considered for download and installation (which can run arbitrary code), or by the site that serves such packages? I don’t see how this can be exploited without the attacker being able to do much more damage; a fix will mainly silence automated checkers. https://github.com/python/cpython/issues/102202
Most ReDOS vulnerabilities are self-attacks, meaning, not a vulnerability. https://github.com/sarbbottam/eslint-find-rules/issues/349
Pretty much all complex regexes are vulnerable to ReDos. I’ve migrated some more important ones to use a token parsing approach instead of regexes. https://github.com/nodemailer/mailparser/issues/378
When evaluating whether something is a vulnerability, you have to look at the attack vector and the respective cost of upgrading. https://github.com/facebook/docusaurus/issues/10491
Anyone using any utils with inputs of arbitrary length runs a performance risk. Even built-ins aren’t immune. https://github.com/bestiejs/platform.js/issues/139
D.5. Analysis of Dependabot Pull Requests for ReDoS
To analyze the handling of ReDoS vulnerabilities in Dependabot pull requests (GitHub, 2025), we used the same methodology as described for GitHub issues. Specifically, we utilized the GitHub REST API to search for Dependabot-generated pull requests that addressed ReDoS vulnerabilities by querying for redos in the title, body, or comments. The dataset spans from June 2020 to January 2025 and includes 1.03 million pull requests with associated comments, ensuring a comprehensive representation of developer engagement with these security updates.
Figure 13. Merge rates of Dependabot pull requests for ReDoS vulnerabilities (June 2020–January 2025), with an average rate of 51.53%.
Figure 13 shows the results of our analysis. While the merge rate for Dependabot pull requests addressing ReDoS vulnerabilities exhibits fluctuations over time, the average merge rate stands at 51.53%. This rate is notably lower than the average merge rate for other types of security pull requests (Alfadel et al., 2021), highlighting that developers often deprioritize ReDoS-related fixes. Reasons for this may include a perception that the vulnerabilities are not critical or relevant, as well as concerns about the complexity or potential disruptions caused by merging these updates.